OpenStudy (anonymous):

what is the solution of the system? Use elimination 5x+8y=-29 7x-2y=-67

OpenStudy (mathstudent55):

Look at the 8y in the first equation and at the -2y in the second equation. If you add the equations the way they are now, nothing will eliminate, but if you multiply the entire second equation by 4, the y term will become -8y, and when added to the 8y of the first equation will be eliminated.

OpenStudy (anonymous):

wait im confused where did 4 come from?

OpenStudy (mathstudent55):

You need to do something so that when you add the 2 equations together one of the variables will add up to zero.

OpenStudy (anonymous):

5x+8y=-29 (Equation A) 7x-2y=-67 (Equation B) ----------- We will be eliminating y first to solve x and then going back to plug in find y. 5x+8y=-29 (Equation A) - leaving it at +8y 7x-2y=-67 (Equation B) 0 multiply all by 4 so we can get -8y 7x*4-2y*4=-67*4 >>> 28x-8y=-268 ----------- We will now subtract downward like you see in the equations below. (A) 5x+8y=-29 (B) 28x-8y=-268 >33x+0y=-297 >33x=-297 >x=-297/33=-9 ----------- Pick either one of the equation (A) or (B) I will pick (A) where 5x+8y=-29 and now we know x=-9 5(-9)+8y=-29 -45+8y=-29 8y=-29+45 8y=16 y=16/8=2 ------------ x=-9 y=2

OpenStudy (mathstudent55):

Since in the frist equation there is 8y, and in the second equation there is -8y, if you multiply the second equation by 4, the -2y becomes -8y which when added to the 8y of the first equation will eliminate the y variable.